OpenStudy (anonymous):

4x^(1/2)-9x^(1/4)+4=0

OpenStudy (anonymous):

\[4x ^{(1/2)}-9x ^{(1/4)}+4 = 0 \]

OpenStudy (anonymous):

\[ u =x ^{ \frac{ 1 }{4}} \]

OpenStudy (anonymous):

wouldn't that be -9x^(1/4)

OpenStudy (anonymous):

im sorry 'u'

OpenStudy (anonymous):

\[4u^{2} -9u +4=0 \]